OpenStudy (anonymous):

Determine the zeros of f(x) = x3 – 3x2 – 16x + 48. (1 point)

OpenStudy (anonymous):

you have lots of possibilities, however 1, -1, 2 and -2 don't work. try 3

OpenStudy (anonymous):

solve by factoring and making f(x) 0 0=x^2-3x^2-16x+48 take out x^2 frm first 2 terms and -16 from last 2 0=x^2(x-3)-16(x-3) combine the x-3 to make 0=(x^2-16)(x-3) factor the perfect square 0=(x-4)(x+4)(x-3) make equal to 0 x-4=0 x+4=0 x-3=0 solve x=4 x=-4 x=3 zeros are (4,0)(-4,0)(3,0)

OpenStudy (anonymous):

and although this will not help you on a quiz or a test, if i wanted to know what the zeros of a cubic polynomial were, i would just ask http://www.wolframalpha.com/input/?i=x3+%E2%80%93+3x2+%E2%80%93+16x+%2B+48

OpenStudy (anonymous):

thank you javier34

OpenStudy (anonymous):

ok